# RedCodeAgent: Automatic Red-teaming Agent against Diverse Code Agents

### v1

@@ 1 Introduction
Large Language Model (LLM)-based code agents are increasingly adopted as assistants to simplify complex coding workflows by generating, refining, and executing code. These agents, often running in information-rich sensitive environments, integrate external tools such as the Python interpreter (Zheng et al., 2024; Wang et al., 2024; Yao et al., 2023) to interact with system environments. As LLM-powered code agents rapidly evolve, their expanding capabilities create new opportunities for automation and problem-solving. However, these agents could also generate and execute buggy or risky code due to security-unaware or even adversarially injected instructions. Such risks can lead to system vulnerabilities, unintended operations, or data breaches (Ruan et al., 2024; Guo et al., 2024), highlighting the need for robust safeguards for code agents.
Traditional red-teaming methods, such as static safety benchmarks and manually designed red-teaming (i.e., jailbreaking) tools, have their own limitations and leave many potential vulnerabilities unexplored. Static benchmarks (Guo et al., 2024; Bhatt et al., 2024; Ruan et al., 2024), while useful for providing baseline safety assessments, are inherently limited in their ability to cover the broad range of users’ boundary behaviors that code agents might encounter in real-world environments (e.g., an adversary might attempt different combinations of attack methods based on the instance from the benchmark). Moreover, manually designed red-teaming algorithms (Zou et al., 2023; Liao & Sun, 2024; Paulus et al., 2024; Liu et al., 2023b) fall short in keeping pace with the rapid evolution of code agents, as they lack the ability to perform dynamic and effective optimization. Traditional red-teaming algorithms can become outdated as agents learn to evade these known patterns. For example, OpenCodeInterpreter (Zheng et al., 2024) incorporates guardrail code with blacklisted commands (e.g., os.remove) into its agent framework to mitigate some found risks, but it can actually be bypassed through alternative implementations (e.g., path.unlink from pathlib). Additionally, in the context of code, to achieve successful red-teaming, it is not enough for the target code agent to merely avoid rejecting the harmful request as in general safety tasks (Zou et al., 2023); the target code agent must generate and execute correct code that performs intended risky functions. Therefore, evaluating execution results beyond static code analysis is also crucial.
@@
We summarize our technical contributions below: 1) We introduce RedCodeAgent, a novel and automated red-teaming agent for evaluating code agents. RedCodeAgent is equipped with an adaptive memory module and a comprehensive toolbox that includes both general-purpose and code-specific red-teaming tools. 2) We develop dedicated simulated environments to assess the execution outcomes of target code agents, avoiding the potential biases introduced by LLM-based evaluators. 3) We conduct a broad evaluation of RedCodeAgent across a variety of security risks – including code generation for malicious applications and Common Weakness Enumeration (CWE) vulnerabilities – spanning multiple programming languages such as Python, C, C++, and Java. 4) We demonstrate the effectiveness of RedCodeAgent, which achieves significantly higher attack success rates and lower rejection rates compared to state-of-the-art LLM jailbreak methods across diverse code agents, including OpenCodeInterpreter (Zheng et al., 2024), ReAct (Liu et al., 2023a), the multi-agent framework MetaGPT (Hong et al., 2024), and commercial agents such as Cursor (Cursor., 2024) and Codeium (Codeium., 2024). 5) We show that RedCodeAgent is both efficient and generalizable, which maintains comparable runtime to a single jailbreak method, while dynamically adapting tool usage based on the risk scenario and red-teaming difficulty. 6) We uncover several notable insights, including the most common vulnerabilities across different agents, variation in red-teaming difficulty across goals, the weaknesses of different code agents, and the frequently triggered attack tools. In addition, we find RedCodeAgent can uncover new vulnerabilities, which other baselines fail to identify.
@@ 2 Related Work
LLM Agent. LLM agents (Yao et al., 2023; Xi et al., 2023), with large language models (LLMs) as their core, implement tasks by interacting with the environment. These agents are often equipped with a memory module, enabling knowledge-based reasoning to handle various tasks within their application domains (Lewis et al., 2020). LLM agents have been deployed for a variety of tasks, such as code generation and execution (Zheng et al., 2024; Wang et al., 2024), as well as red teaming. For example, Xu et al. (2024) proposed a general agent framework for jailbreaking (static) LLMs, while Fang et al. (2024) demonstrated agents can exploit one-day vulnerabilities. However, none of the red teaming work targets code agents, which involves additional complexity in code generation and execution tasks.
@@
Agent Safety. Existing agent safety benchmarks, such as ToolEmu (Ruan et al., 2024), R-judge (Yuan et al., 2024), AgentMonitor (Naihin et al., 2023) and HAICOSYSTEM (Zhou et al., 2024), focus on providing datasets of risky interaction records and utilize LLMs as judges to identify safety risks within the provided records. In contrast, our goal is to conduct direct red-teaming against given code agents. Recently, Guo et al. (2024) introduced a safety benchmark specifically designed for code agents. However, this benchmark relies heavily on extensive human labor, and as agents evolve rapidly, static benchmarks can quickly become outdated. Current red-teaming strategies, such as memory poisoning attacks on agents (Chen et al., 2024b), often lack automation and are not comprehensive. In contrast, our proposed RedCodeAgent, offers a fully automated and adaptive red-teaming methodology, addressing the shortcomings of existing strategies.
@@
Safety of Code LLMs. Existing benchmarks (Bhatt et al., 2023; 2024; Peng et al., 2025; Pa et al., 2023; Pearce et al., 2022; Yang et al., 2024; Hajipour et al., 2024) have revealed that code LLMs may generate unsafe code snippets. Code agents, however, differ from traditional code LLMs in several key aspects. Code agents are more complex, often featuring multi-round self-refinement (Zheng et al., 2024), and can directly interact with and modify the user’s environment. Unlike prior work that primarily evaluates risks in static code generated by LLMs, our focus extends to the safety implications of the actions agents take in diverse execution environments. To ensure reliable evaluation, our design includes a specialized sandbox for code execution and carefully tailored test cases. While our approach is designed for code agents, it can also generalize to traditional code LLMs, offering a flexible framework. Existing code LLM red teaming methods aim to elicit risky code from code LLMs. While our work focuses on adversarial attacks on the code generation task under a black-box setting, where the input consists of natural language and the output is code, prior work has targeted different tasks. CodeAttack (Jha & Reddy, 2023) focuses on code translation, code summarization, and code completion tasks. INSEC (Jenko et al., 2024) focuses on code completion, where the input is part of the code. SVEN (He & Vechev, 2023) operates in the white-box setting and proposes methods to train models to generate safe or unsafe code. Few attacks have been directed specifically at the code generation task,Cotroneo et al. (2024); Aghakhani et al. (2024) introduce vulnerabilities by adding malicious code to the training dataset rather than directly attacking deployed models. These contrasts highlight the novelty of RedCodeAgent, which explores an underexamined aspect of adversarial attacks on black-box code generation tasks.
@@ 3.1 Overview of RedCodeAgent
Threat Model. RedCodeAgent aims to perform automated red-teaming penetration tests to evaluate the security of target code agents. We consider potential adversaries who may provide risky instructions to mislead target code agents to generate or execute risky code. We assume code agents execute code without additional human intervention. This is a practical scenario, as even advanced safety confirmation steps might be bypassed under inattentive supervision, leading to potential vulnerabilities, as discussed by prior work (Liao et al., 2024; 2025).
@@ 3.3.1 Code Substitution
During experiments, we observed that due to the sensitive nature of some red-teaming functionalities, the Code Substitution LLM occasionally refuses to fulfill certain requests. To address this issue, we selected several candidate LLMs for Code Substitution: CodeQwen1.5-7B-Chat (Bai et al., 2023), Meta-Llama-3-8B-Instruct (AI@Meta, 2024), and gpt-35-turbo (Achiam et al., 2023). When one LLM rejects a request, another LLM is selected to obtain the required code implementation.
@@ 3.3.2 General Jailbreak Tools
Existing studies (Jin et al., 2024; Yi et al., 2024) have demonstrated the effectiveness of jailbreak attacks in general NLP tasks, showing that such methods can reduce the likelihood of the target LLM (or agent) rejecting a given request. In our work, we include multiple representative jailbreak approaches to ensure comprehensive coverage: gradient-based attacks such as GCG (Zou et al., 2023); learning-based attacks including Advprompter (Paulus et al., 2024) and AmpleGCG (Liao & Sun, 2024); and evolutionary-based attacks such as AutoDAN (Liu et al., 2023b). Given the scalability of RedCodeAgent, users can easily extend the framework with additional jailbreak techniques. In §˜D.7, we also introduce additional template-based and role-play-based attacks.
@@ 3.4 Evaluation Module
Risk Scenarios. We use the risk scenarios (i.e., test cases) from RedCode-Exec (see details in Fig.˜6), RedCode-Gen (Guo et al., 2024) and RMCbench (Chen et al., 2024a) as input for RedCodeAgent. The test cases in RedCode-Exec are prompts that guide the target code agent to generate unsafe code and possibly execute risky code. In RedCode-Exec, 14 code-related risk scenarios were collected from existing agent safety benchmarks (Yuan et al., 2024), and 13 risk scenarios were derived from MITRE’s Common Weakness Enumeration (CWE) (The MITRE Corporation, 2024). The above risk scenarios span 8 domains, including operating systems, file systems, and others. In RedCode-Exec, each risk scenario contains 30 natural language test cases, resulting in test cases. The test cases in RedCode-Gen are function docstring instructions derived from eight malware families that guide the target code agent to generate malware with explicit malicious intent and the potential to cause harm. There is a total of 160 test cases in RedCode-Gen. The test cases in RMCbench evaluates the ability of LLMs to resist generating malicious code, we use all the 182 text-to-code test cases in RMCbench for our experiments. Additionally, we discussed prompt injection attack in Tb.˜9 and the SQL injection risk scenario in §˜D.11.
@@
Simulational Evaluation Environment. In the context of code, where the interaction involves a target code agent executing code, an execution environment and concrete evaluation scripts corresponding to each risk description can be implemented to ensure rigorous evaluation. For example, in tasks such as deleting sensitive files, after the code agent receives the request and potentially executes the code in a sandbox, we can directly examine the file system in the sandbox to determine whether the risk description was realized. Also, sandbox isolation allows us to test potentially dangerous code without compromising system security. Following these criteria, we build a Docker (Merkel et al., 2014) container where we prepare the necessary environment setups and resources to ensure successful code execution. We also create evaluation scripts corresponding to each risk description to assess the outcome of the code execution. The evaluation result falls into one of the following categories: (1) Rejection: The user query is rejected. (2) Execution Failure: The threat functionality is not achieved. (3) Attack Success: The threat functionality is successfully achieved (i.e., code execution succeeds and brings the corresponding outcome). An example of such evaluation is shown in §˜E.1. The evaluation module provides more detailed feedback beyond merely the keywords like “Attack Success". For example, in the case of deleting sensitive files, the evaluation module would explain the reason for this evaluation result, such as attack failure due to the target file still existing. We find this explanation to be a helpful reminder for RedCodeAgent, ensuring that RedCodeAgent does not forget the attack objective. The response of the code agent, evaluation result, and the reason for this evaluation result serve as feedback to RedCodeAgent, enabling it to further optimize its attack strategies.
@@ 4 Experimental Results
Baselines and Metrics. We consider 4 state-of-the-art jailbreak methods GCG (Zou et al., 2023), AmpleGCG (Liao & Sun, 2024), Advprompter (Paulus et al., 2024), and AutoDAN (Liu et al., 2023b) as our baselines, which demonstrate strong jailbreak performance in general safety tasks. For these baselines, we applied their corresponding optimization methods to optimize the static test cases and used the optimized prompts as test cases for the code agent. We also consider No Jailbreak as another baseline, which refers to directly using static test cases (from the RedCode-Exec or RedCode-Gen dataset) as input to the target code agent. Three metrics are reported in the main paper: attack success rate (ASR), rejection rate (RR), and time cost. We also compare the perceived stealthiness of the prompt optimized by different methods in §˜D.12. We consider the following targeted code agents: OpenCodeInterpreter (Zheng et al., 2024), ReAct (Liu et al., 2023a), the multi-agent framework MetaGPT Hong et al. (2024), and commercial agents such as Cursor (Cursor., 2024) and Codeium (Codeium., 2024).
RedCodeAgent Setup. RedCodeAgent is built on LangChain framework (Topsakal & Akinci, 2023), with GPT-4o-mini (Achiam et al., 2023) as its base LLM. We follow the memory structure design outlined in §˜3.2, and the tools provided to RedCodeAgent adhere to the setup described in §˜3.3. We set the max_iterations to 35 to control the total number of iterations. For the memory search, we use sentence-transformers/paraphrase-MiniLM-L6-v2 (Reimers & Gurevych, 2019) as our embedding model. We set top , meaning RedCodeAgent receives the three most similar successful attack experiences (if fewer than are available in the memory, all successful entries are provided). The penalty factor . RedCodeAgent dynamically accumulates successful experiences by starting with an empty memory and executing test cases sequentially. After each case, successful experiences are stored in memory, allowing the agent to leverage prior knowledge when tackling subsequent cases. The details about the mechanism of memory accumulation are described in §˜F.1.4. Other detailed experimental settings are provided in §˜F.

### v2

@@ 1 Introduction
Large Language Model (LLM)-based code agents are increasingly adopted as assistants to simplify complex coding workflows by generating, refining, and executing code. These agents, often running in information-rich sensitive environments, integrate external tools such as the Python interpreter (zheng2024opencodeinterpreter; wang2024executable; yao2023react) to interact with system environments. As LLM-powered code agents rapidly evolve, their expanding capabilities create new opportunities for automation and problem-solving. However, these agents could also generate and execute buggy or risky code due to security-unaware or even adversarially injected instructions. Such risks can lead to system vulnerabilities, unintended operations, or data breaches (ruan2024toolemu; guo2024redcode), highlighting the need for robust safeguards for code agents.
Traditional red-teaming methods, such as static safety benchmarks and manually designed red-teaming (i.e., jailbreaking) tools, have their own limitations and leave many potential vulnerabilities unexplored. Static benchmarks (guo2024redcode; bhatt2024cyberseceval; ruan2024toolemu), while useful for providing baseline safety assessments, are inherently limited in their ability to cover the broad range of users’ boundary behaviors that code agents might encounter in real-world environments (e.g., an adversary might attempt different combinations of attack methods based on the instance from the benchmark). Moreover, manually designed red-teaming algorithms (zou2023universal; liao2024amplegcg; paulus2024advprompter; liu2023autodan) fall short in keeping pace with the rapid evolution of code agents, as they lack the ability to perform dynamic and effective optimization. Traditional red-teaming algorithms can become outdated as agents learn to evade these known patterns. For example, OpenCodeInterpreter (zheng2024opencodeinterpreter) incorporates guardrail code with blacklisted commands (e.g., os.remove) into its agent framework to mitigate some found risks, but it can actually be bypassed through alternative implementations (e.g., path.unlink from pathlib). Additionally, in the context of code, to achieve successful red-teaming, it is not enough for the target code agent to merely avoid rejecting the harmful request as in general safety tasks (zou2023universal); the target code agent must generate and execute correct code that performs intended risky functions. Therefore, evaluating execution results beyond static code analysis is also crucial.
@@
We summarize our technical contributions below: 1) We introduce RedCodeAgent, a novel and automated red-teaming agent for evaluating code agents. RedCodeAgent is equipped with an adaptive memory module and a comprehensive toolbox that includes both general-purpose and code-specific red-teaming tools. 2) We develop dedicated simulated environments to assess the execution outcomes of target code agents, avoiding the potential biases introduced by LLM-based evaluators. 3) We conduct a broad evaluation of RedCodeAgent across a variety of security risks – including code generation for malicious applications and Common Weakness Enumeration (CWE) vulnerabilities – spanning multiple programming languages such as Python, C, C++, and Java. 4) We demonstrate the effectiveness of RedCodeAgent, which achieves significantly higher attack success rates and lower rejection rates compared to state-of-the-art LLM jailbreak methods across diverse code agents, including OpenCodeInterpreter (zheng2024opencodeinterpreter), ReAct (liu2023agentbench), the multi-agent framework MetaGPT (hong2024metagpt), and commercial agents such as Cursor (cursor2024) and Codeium (codeium2024). 5) We show that RedCodeAgent is both efficient and generalizable, which maintains comparable runtime to a single jailbreak method, while dynamically adapting tool usage based on the risk scenario and red-teaming difficulty. 6) We uncover several notable insights, including the most common vulnerabilities across different agents, variation in red-teaming difficulty across goals, the weaknesses of different code agents, and the frequently triggered attack tools. In addition, we find RedCodeAgent can uncover new vulnerabilities, which other baselines fail to identify.
@@ 2 Related Work
LLM Agent. LLM agents (yao2023react; xi2023rise), with large language models (LLMs) as their core, implement tasks by interacting with the environment. These agents are often equipped with a memory module, enabling knowledge-based reasoning to handle various tasks within their application domains (lewis2020retrieval). LLM agents have been deployed for a variety of tasks, such as code generation and execution (zheng2024opencodeinterpreter; wang2024executable), as well as red teaming. For example, xu2024redagent proposed a general agent framework for jailbreaking (static) LLMs, while fang2024llm demonstrated agents can exploit one-day vulnerabilities. However, none of the red teaming work targets code agents, which involves additional complexity in code generation and execution tasks.
@@
Agent Safety. Existing agent safety benchmarks, such as ToolEmu (ruan2024toolemu), R-judge (yuan2024rjudge), AgentMonitor (naihin2023testing) and HAICOSYSTEM (zhou2024haicosystemecosystemsandboxingsafety), focus on providing datasets of risky interaction records and utilize LLMs as judges to identify safety risks within the provided records. In contrast, our goal is to conduct direct red-teaming against given code agents. Recently, guo2024redcode introduced a safety benchmark specifically designed for code agents. However, this benchmark relies heavily on extensive human labor, and as agents evolve rapidly, static benchmarks can quickly become outdated. Current red-teaming strategies, such as memory poisoning attacks on agents (chen2024agentpoison), often lack automation and are not comprehensive. In contrast, our proposed RedCodeAgent, offers a fully automated and adaptive red-teaming methodology, addressing the shortcomings of existing strategies.
@@
Safety of Code LLMs. Existing benchmarks (bhatt2023purple; bhatt2024cyberseceval; peng2025cweval; PaPa2023AnAD; pearce2022asleep; yang2024sweagent; hajipour2024codelmsec) have revealed that code LLMs may generate unsafe code snippets. Code agents, however, differ from traditional code LLMs in several key aspects. Code agents are more complex, often featuring multi-round self-refinement (zheng2024opencodeinterpreter), and can directly interact with and modify the user’s environment. Unlike prior work that primarily evaluates risks in static code generated by LLMs, our focus extends to the safety implications of the actions agents take in diverse execution environments. To ensure reliable evaluation, our design includes a specialized sandbox for code execution and carefully tailored test cases. While our approach is designed for code agents, it can also generalize to traditional code LLMs, offering a flexible framework. Existing code LLM red teaming methods aim to elicit risky code from code LLMs. While our work focuses on adversarial attacks on the code generation task under a black-box setting, where the input consists of natural language and the output is code, prior work has targeted different tasks. CodeAttack (jha2023codeattack) focuses on code translation, code summarization, and code completion tasks. INSEC (jenko2024practical) focuses on code completion, where the input is part of the code. SVEN (he2023large) operates in the white-box setting and proposes methods to train models to generate safe or unsafe code. Few attacks have been directed specifically at the code generation task, cotroneo2024vulnerabilities and aghakhani2024trojanpuzzle introduce code vulnerability by adding malicious code to the training dataset rather than directly attacking deployed models. These contrasts highlight the novelty of RedCodeAgent, which explores an underexamined aspect of adversarial attacks on black-box code generation tasks.
@@ 3.1 Overview of RedCodeAgent
Threat Model. RedCodeAgent aims to perform automated red-teaming penetration tests to evaluate the security of target code agents. We consider potential adversaries who may provide risky instructions to mislead target code agents to generate or execute risky code. We assume code agents execute code without additional human intervention. This is a practical scenario, as even advanced safety confirmation steps might be bypassed under inattentive supervision, leading to potential vulnerabilities, as discussed by prior work (liao2024eia; liao2025redteamcua).
@@ 3.3.1 Code Substitution
During experiments, we observed that due to the sensitive nature of some red-teaming functionalities, the Code Substitution LLM occasionally refuses to fulfill certain requests. To address this issue, we selected several candidate LLMs for Code Substitution: CodeQwen1.5-7B-Chat (qwen), Meta-Llama-3-8B-Instruct (llama3modelcard), and gpt-35-turbo (achiam2023gpt). When one LLM rejects a request, another LLM is selected to obtain the required code implementation.
@@ 3.3.2 General Jailbreak Tools
Existing studies (jin2024jailbreakzoo; yi2024jailbreak) have demonstrated the effectiveness of jailbreak attacks in general NLP tasks, showing that such methods can reduce the likelihood of the target LLM (or agent) rejecting a given request. In our work, we include multiple representative jailbreak approaches to ensure comprehensive coverage: gradient-based attacks such as GCG (zou2023universal); learning-based attacks including Advprompter (paulus2024advprompter) and AmpleGCG (liao2024amplegcg); and evolutionary-based attacks such as AutoDAN (liu2023autodan). Given the scalability of RedCodeAgent, users can easily extend the framework with additional jailbreak techniques. In §˜D.7, we also introduce additional template-based and role-play-based attacks.
@@ 3.4 Evaluation Module
Risk Scenarios. We use the risk scenarios (i.e., test cases) from RedCode-Exec (see details in Fig.˜6), RedCode-Gen (guo2024redcode) and RMCbench (chen2024rmcbench) as input for RedCodeAgent. The test cases in RedCode-Exec are prompts that guide the target code agent to generate unsafe code and possibly execute risky code. In RedCode-Exec, 14 code-related risk scenarios were collected from existing agent safety benchmarks (yuan2024rjudge), and 13 risk scenarios were derived from MITRE’s Common Weakness Enumeration (CWE) (cwev414). The above risk scenarios span 8 domains, including operating systems, file systems, and others. In RedCode-Exec, each risk scenario contains 30 natural language test cases, resulting in test cases. The test cases in RedCode-Gen are function docstring instructions derived from eight malware families that guide the target code agent to generate malware with explicit malicious intent and the potential to cause harm. There is a total of 160 test cases in RedCode-Gen. The test cases in RMCbench evaluates the ability of LLMs to resist generating malicious code, we use all the 182 text-to-code test cases in RMCbench for our experiments. Additionally, we discussed prompt injection attack in Tb.˜9 and the SQL injection risk scenario in §˜D.11.
@@
Simulational Evaluation Environment. In the context of code, where the interaction involves a target code agent executing code, an execution environment and concrete evaluation scripts corresponding to each risk description can be implemented to ensure rigorous evaluation. For example, in tasks such as deleting sensitive files, after the code agent receives the request and potentially executes the code in a sandbox, we can directly examine the file system in the sandbox to determine whether the risk description was realized. Also, sandbox isolation allows us to test potentially dangerous code without compromising system security. Following these criteria, we build a Docker (merkel2014docker) container where we prepare the necessary environment setups and resources to ensure successful code execution. We also create evaluation scripts corresponding to each risk description to assess the outcome of the code execution. The evaluation result falls into one of the following categories: (1) Rejection: The user query is rejected. (2) Execution Failure: The threat functionality is not achieved. (3) Attack Success: The threat functionality is successfully achieved (i.e., code execution succeeds and brings the corresponding outcome). An example of such evaluation is shown in §˜E.1. The evaluation module provides more detailed feedback beyond merely the keywords like “Attack Success". For example, in the case of deleting sensitive files, the evaluation module would explain the reason for this evaluation result, such as attack failure due to the target file still existing. We find this explanation to be a helpful reminder for RedCodeAgent, ensuring that RedCodeAgent does not forget the attack objective. The response of the code agent, evaluation result, and the reason for this evaluation result serve as feedback to RedCodeAgent, enabling it to further optimize its attack strategies.
@@ 4 Experimental Results
Baselines and Metrics. We consider 4 state-of-the-art jailbreak methods GCG (zou2023universal), AmpleGCG (liao2024amplegcg), Advprompter (paulus2024advprompter), and AutoDAN (liu2023autodan) as our baselines, which demonstrate strong jailbreak performance in general safety tasks. For these baselines, we applied their corresponding optimization methods to optimize the static test cases and used the optimized prompts as test cases for the code agent. We also consider No Jailbreak as another baseline, which refers to directly using static test cases (from the RedCode-Exec or RedCode-Gen dataset) as input to the target code agent. Three metrics are reported in the main paper: attack success rate (ASR), rejection rate (RR), and time cost. We also compare the perceived stealthiness of the prompt optimized by different methods in §˜D.12. We consider the following targeted code agents: OpenCodeInterpreter (zheng2024opencodeinterpreter), ReAct (liu2023agentbench), the multi-agent framework MetaGPT (hong2024metagpt), and commercial agents such as Cursor (cursor2024) and Codeium (codeium2024).
RedCodeAgent Setup. RedCodeAgent is built on LangChain framework (topsakal2023creating), with GPT-4o-mini (achiam2023gpt) as its base LLM. We follow the memory structure design outlined in §˜3.2, and the tools provided to RedCodeAgent adhere to the setup described in §˜3.3. We set the max_iterations to 35 to control the total number of iterations. For the memory search, we use sentence-transformers/paraphrase-MiniLM-L6-v2 (reimers-2019-sentence-bert) as our embedding model. We set top , meaning RedCodeAgent receives the three most similar successful attack experiences (if fewer than are available in the memory, all successful entries are provided). The penalty factor . RedCodeAgent dynamically accumulates successful experiences by starting with an empty memory and executing test cases sequentially. After each case, successful experiences are stored in memory, allowing the agent to leverage prior knowledge when tackling subsequent cases. The details about the mechanism of memory accumulation are described in §˜F.1.4. Other detailed experimental settings are provided in §˜F.